Solve the system algebraically:
x-2y=1
y= sq root of (x-1)

Answer by stanbon(75887)

Solve the system algebraically:
x-2y=1
x = 2y+1
y= sq root of (x-1)
-----------------------
Substitute for "x" and solve for "y":
y = sqrt(2y+1-1)
------
y = sqrt(2y)
y^2 = 4y
y^2-4y = 0
y(y-4) = 0
y = 0 or y = 4
----
If y = 0, x = 2y+1 = 1
If y = 4, x = 2y+1 = 9
===========================
Cheers,
Stan H.